OpenStudy (anonymous):

if f(x) = 8x^2 then what does f(x+h) = ?

OpenStudy (anonymous):

Well, the difference between f(x) and f(x+h) is simply that x is replaced with x + h. So in your function f(x) = 8x^2, replace x with x + h. Thats all ya gotta do :)

OpenStudy (anonymous):

oh goodness thanks! i was attempting factoring and etc. lol

OpenStudy (anonymous):

Yeah xD So another example just to show the idea. If I have f(x) = 7x^3 + 8x^2 - 3 and I ask to find f(x+y+z). Well, all that happened was x was replaced with x+y+z. So thats all you would do is make that replacement. So f(x+y+z) = 7(x+y+z)^3 + 8(x+y+z)^2 - 3. Whatever is inside the f(__) you replace every x with.

OpenStudy (anonymous):

ive attemted in it but i am being told it is still wrong here is what i did. f(x) = 8x2 -->f(x+h)= 8 (x+h)^2 = 8x^2+16hx+h^2 what did i do wrong?

OpenStudy (anonymous):

You didnt even need to foil it out. f(x+h) = 8(x+h)^2 is perfectly fine if it were up to me.

OpenStudy (anonymous):

I did that too and still told it was wrong

OpenStudy (anonymous):

If you actually had to foil it out, you'd have 8(x^2+2xh+h^2) = 8x^2 + 16xh + 8h^2

OpenStudy (anonymous):

For the function, find and simplify f(x + h). (Expand your answer completely.) f(x) = 8x2 f(x + h) =

OpenStudy (anonymous):

Okay, expanded, so they want the 8x^2 + 16xh + 8h^2 answer.

OpenStudy (anonymous):

oh i am an idiot I did not put 8h^2 i wrote in h^2 when I foiled it hahahaha

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

Yep, np :)